# PANORAMIC: Discovery of an Ultra-Massive Grand-Design Spiral Galaxy at

### v1

@@ 3.4 The location on the main sequence
Given the red, quiescent-like bulge, we investigate the location of our galaxy with respect to the star-formation main sequence (SFMS) at . While previous ultra-massive galaxies are found to be in a rapid star-formation phase, Zhúlóng’s star-formation rate is significantly low. For example, the three sources from Xiao et al. (2024) S1, S2, and S3 are ultra-massive, dusty star-forming, sub-millimeter galaxies. They have SFRs of 1000 yr-1, large dust contents ( mag and detected by SCUBA2 observations), half-light radii of kpc. Due to heavy dust obscuration, they are barely detectable in the optical wavelength bands, and hence they are also called optically dark galaxies (or HST-dark galaxies; e.g., Wang et al., 2019; Xiao et al., 2023; Williams et al., 2024a). By contrast, as shown in Fig. 5, the SFR of Zhúlóng is 66 yr-1 (also in Table 1), more than 10 times lower than S1, S2, and S3. It is not detected in the deep ALMA observations at 1.2mm (3 upper limit of 0.6 mJy beam-1) and other FIR observations, consistent with its low measured SFR (see Sect. 2.2). With these properties, the galaxy is about 1 dex below the main sequence as characterized by Schreiber et al. (2015), and 0.5 dex below the relation of Popesso et al. (2023).
@@ 4.1 An outlier among massive galaxies at
However, in light of its low integrated sSFR ( sSFR yr-1; 0.5 dex below the SFMS at these redshifts; see Fig. 5), such a large size is quite remarkable. Its red rest-frame U-V vs V-J colors in the core (see Section 3.2) are consistent with quiescent systems, perhaps making a comparison to that population more appropriate. Typically, UVJ-red galaxies follow a steeper mass-size relation leading to much smaller sizes at fixed stellar mass than SFGs (e.g., van der Wel et al., 2014; Ito et al., 2024; Ji et al., 2024b). Quiescent galaxies with similarly high mass (10.3-11) at comparable redshifts (; e.g., Carnall et al., 2023; de Graaff et al., 2024; Weibel et al., 2024a; Ji et al., 2024a; Ito et al., 2024), have Re kpc, a factor of smaller than what we find here (using our single-Sérsic measurement for Zhúlóng, for consistency with other studies). Unfortunately, sample sizes of comparable galaxies at at remain small, regardless of sSFR, and precludes a robust conclusion in comparison to the population scatter.
@@ 5 Conclusions
In this paper, we present the serendipitous discovery of a new ultra-massive galaxy identified in the pure parallel imaging survey PANORAMIC (in the COSMOS field). It has well-constrained by a strong Balmer/4000 Å break and consistent with the neighboring star-forming clump (Fig. 1). After combining the JWST/NIRCam, HST, and ALMA data, we find the following extremely intriguing properties:
@@
Altogether, Zhúlóng at is a red, ultra-massive, grand-design spiral galaxy with a large stellar size. Overall the galaxy sits below the main sequence, suggesting the star formation activity is ramping down. It contains a quiescent bulge + large star-forming disk + spiral arms. For a galaxy at , having any of these relevant properties would be enough to make the galaxy special. The presence of all of these properties makes Zhúlóng very exceptional, indicating the rapid formation and morphological evolution of massive galaxies in the early universe. With future JWST IFU observations, combined with higher-resolution and deeper ALMA data, this galaxy will provide a unique window to unveil what triggers the rapid mass assembly together with the early disk formation.
@@ Acknowledgements.
We thank Boris Sindhu Kalita, Francoise Combes, and Maximilien Franco for valuable discussions and suggestions that improved this paper. This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 2514. Support for program JWST-GO-2514 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2023.1.00180.L. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. The work of CCW is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. PD acknowledges support from the NWO grant 016.VIDI.189.162 (“ODIN”) and warmly thanks the European Commission’s and University of Groningen’s CO-FUND Rosalind Franklin program. AH acknowledges support by the VILLUM FONDEN under grant 37459. MVM acknowledges support from the National Science Foundation via AAG grant 2205519 and the Wisconsin Alumni Research Foundation via grant MSN251397. CCW, AC, ZJ and KEW acknowledge funding from program JWST-GO-2514 that was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. ZJ also acknowledges funding from JWST/NIRCam contract to the University of Arizona NAS5-02015. I.L. acknowledges support from Australian Research Council Future Fellowship FT220100798. SL acknowledges support by the Science and Technology Facilities Council (STFC) and by the UKRI Frontier Research grant RISEandFALL. LB acknowledges funding from STFC through ST/T000473/1 and ST/X001040/1. K.G. and T.N. acknowledge support from Australian Research Council Laureate Fellowship FL180100060.References
@@
- Williams et al. (2024a) Williams, C. C., Alberts, S., Ji, Z., et al. 2024a, ApJ, 968, 34
- Williams et al. (2024b) Williams, C. C., Oesch, P. A., Weibel, A., et al. 2024b, arXiv e-prints, arXiv:2410.01875
@@
Appendix A Morphological modeling of Zhúlóng.

### v2

@@ 3.4 The location on the main sequence
Given the red, quiescent-like bulge, we investigate the location of our galaxy with respect to the star-formation main sequence (SFMS) at . While previous ultra-massive galaxies are found to be in a rapid star-formation phase, Zhúlóng’s star-formation rate is significantly low. For example, the three sources from Xiao et al. (2024) S1, S2, and S3 are ultra-massive, dusty star-forming, sub-millimeter galaxies. They have SFRs of 1000 yr-1, large dust contents ( mag and detected by SCUBA2 observations), half-light radii of kpc. Due to heavy dust obscuration, they are barely detectable in the optical wavelength bands, and hence they are also called optically dark galaxies (or HST-dark galaxies; e.g., Wang et al., 2019; Williams et al., 2019; Xiao et al., 2023; Williams et al., 2024). By contrast, as shown in Fig. 5, the SFR of Zhúlóng is 66 yr-1 (also in Table 1), more than 10 times lower than S1, S2, and S3. It is not detected in the deep ALMA observations at 1.2mm (3 upper limit of 0.6 mJy beam-1) and other FIR observations, consistent with its low measured SFR (see Sect. 2.2). With these properties, the galaxy is about 1 dex below the main sequence as characterized by Schreiber et al. (2015), and 0.5 dex below the relation of Popesso et al. (2023).
@@ 4.1 An outlier among massive galaxies at
However, in light of its low integrated sSFR ( sSFR yr-1; 0.5 dex below the SFMS at these redshifts; see Fig. 5), such a large size is quite remarkable. Its red rest-frame U-V vs V-J colors in the core (see Section 3.2) are consistent with quiescent systems, perhaps making a comparison to that population more appropriate. Typically, UVJ-red galaxies follow a steeper mass-size relation leading to much smaller sizes at fixed stellar mass than SFGs (e.g., van der Wel et al., 2014; Ito et al., 2024; Ji et al., 2024b). Quiescent galaxies with similarly high mass (10.3-11) at comparable redshifts (; e.g., Carnall et al., 2023; de Graaff et al., 2024; Weibel et al., 2024a; Ji et al., 2024a; Ito et al., 2024), have Re kpc, a factor of smaller than what we find here (using our single-Sérsic measurement for Zhúlóng, for consistency with other studies). Unfortunately, sample sizes of comparable galaxies at at remain small, regardless of sSFR, and preclude a robust conclusion in comparison to the population scatter.
@@ 5 Conclusions
In this paper, we present the serendipitous discovery of the most distant ultra-massive spiral galaxy candidate identified in the pure parallel imaging survey PANORAMIC (in the COSMOS field). It has well-constrained by a strong Balmer/4000 Å break and consistent with the neighboring star-forming clump (Fig. 1). After combining the JWST/NIRCam, HST, and ALMA data, we find the following extremely intriguing properties:
@@
Altogether, Zhúlóng at is potentially the highest-redshift, red, ultra-massive, grand-design spiral galaxy with a large stellar size. Overall, the galaxy sits below the main sequence, suggesting the star formation activity is ramping down. It contains a quiescent bulge + large star-forming disk + spiral arms. For a galaxy at , having any of these relevant properties would be enough to make the galaxy special. The presence of all of these properties makes Zhúlóng very exceptional, indicating the rapid formation and morphological evolution of massive galaxies in the early universe. With future JWST IFU observations, combined with higher-resolution and deeper ALMA data, this galaxy will provide a unique window to unveil what triggers the rapid mass assembly together with the early disk formation.
@@ Acknowledgements.
We are very grateful to the anonymous referee for instructive comments, which helped to improve the overall quality and strengthened the analyses of this work. We thank Boris Sindhu Kalita, Francoise Combes, and Maximilien Franco for valuable discussions and suggestions that improved this paper. This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program 2514. Support for program JWST-GO-2514 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2023.1.00180.L. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), NSTC and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. The work of CCW is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. PD acknowledges support from the NWO grant 016.VIDI.189.162 (“ODIN”) and warmly thanks the European Commission’s and University of Groningen’s CO-FUND Rosalind Franklin program. AH acknowledges support by the VILLUM FONDEN under grant 37459. MVM acknowledges support from the National Science Foundation via AAG grant 2205519 and the Wisconsin Alumni Research Foundation via grant MSN251397. CCW, AC, ZJ and KEW acknowledge funding from program JWST-GO-2514 that was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. ZJ also acknowledges funding from JWST/NIRCam contract to the University of Arizona NAS5-02015. I.L. acknowledges support from Australian Research Council Future Fellowship FT220100798. SL acknowledges support by the Science and Technology Facilities Council (STFC) and by the UKRI Frontier Research grant RISEandFALL. LB acknowledges funding from STFC through ST/T000473/1 and ST/X001040/1. K.G. and T.N. acknowledge support from Australian Research Council Laureate Fellowship FL180100060.References
@@
- Williams et al. (2024) Williams, C. C., Alberts, S., Ji, Z., et al. 2024, ApJ, 968, 34
- Williams et al. (2019) Williams, C. C., Labbe, I., Spilker, J., et al. 2019, ApJ, 884, 154
- Williams et al. (2025) Williams, C. C., Oesch, P. A., Weibel, A., et al. 2025, ApJ, 979, 140
@@
Appendix A Robustness of the redshift solution for Zhúlóng.
In Fig. 1, we show the narrow photometric redshift distribution of Zhúlóng, derived using EAZY for the galaxy core (0.16′′ radius aperture). To ensure the robustness of our results, we also perform SED fitting with Bagpipes, allowing the redshift to vary from 1 to 10. We assume different SFHs, including delayed SFH and double-power-law SFH, and the same settings as described in Sect. 3.2. The resulting shows a single strong peak, consistent with the EAZY results. Specifically, the photometric redshift is for delayed SFH (Fig. 6-) and for the double-power-law SFH.
However, when a larger aperture (e.g., 0.7′′ radius) is used, a second redshift solution at emerges (Fig. 6-). To further investigate this, we rerun the SED fitting for the galaxy core using EAZY with the redshift fixed to . The resulting fit, shown in Fig. 6-, has a large chi-square value (), indicating a very poor fit.
To explore the cause of this low-redshift solution with a large aperture, we further inspect the images and find a bright clump (clump 2, thereafter; Fig. 6-) at F606W and F814W from HST/ACS. Clump 2 is located south-east of the clump reported in Fig. 1. The EAZY fitting results for clump 2 show a broad distribution with , suggesting it may be a foreground clump. Thus, the second redshift peak in Fig. 6- is likely due to contamination from clump 2.
Overall, we perform SED fitting on the galaxy core using both EAZY and Bagpipes, and both methods consistently indicate that Zhúlóng is a strong spiral galaxy candidate at . However, we identify the potential presence of foreground structure(s) that could contaminate the redshift measurements of Zhúlóng when larger apertures are used. We note here that clump 2 does not affect our main results, as the global properties of Zhúlóng are derived using the total flux measured within a 0.5′′ radius aperture (with aperture corrections applied), with clump 2 not included. While clump 2 is included in the SED fitting for the annular radius of 0.5 and may influence the fitting results (see Fig. 3), we cannot confirm whether it is a low- clump without spectroscopic observations and thus did not remove it from the analysis. In any case, the key findings of inside-out galaxy growth and a quiescent-like galaxy core embedded in a star-forming stellar disk remain consistent even when considering only the remaining annuli. In addition, it does not affect our identification of spiral features. To confirm whether Zhúlóng is indeed a distant spiral galaxy at , follow-up JWST NIRSpec/IFU observations are necessary.
Appendix B Morphological modeling of Zhúlóng.